# Asymmetric Proximal Policy Optimization: mini-critics boost LLM reasoning

### v1

@@ Abstract
Figure 1: (a): The initial representational ability of the model makes asymmetric PPO possible, i.e., mini critics guide larger actors. By optimizing the ensemble critic system, AsyPPO achieves reliable value estimation while remaining lightweight. (b): Off-policy ratio=8, Report the average accuracy of 6 benchmarks, i.e., AIME 24, AIME 25, MATH-500, OlympiadBench, MinervaMath, and AMC 2023. (c): The average clock time of the training step and the peak GPU memory usage of AsyPPO are significantly lower than those of the classic PPO, remain at the GRPO level.
@@ 1 Introduction
Faced with these challenges, the RL4LLM community has largely sidelined a key element of classical PPO – its critic. GRPO (He et al., 2025), and its variants, including GSPO (Zheng et al., 2025) in the Qwen series and DAPO (Yu et al., 2025), have achieved great success in replacing value functions with group sampling and average-advantage baselines for coarse-grained estimation of advantages. While effective, this paradigmatic shift abandons a key concept of RL: robust state value estimation can naturally mitigate training collapse caused by advantage bias (Wang et al., 2025b, Liu et al., 2024), especially under off-policy settings. This landscape motivates a fundamental reconsideration of architectural assumptions inherited from deep RL [^1] Please refer to Appendix A for detailed related work. * in the author board denotes the equal contribution., prompting the following central question:
@@
To unlock the capabilities of the small critic, we consider critic ensembles to improve its value estimation and policy guidance. However, naive ensembles offer limited benefits for policy learning, as LLM critics start from identical pre-trained checkpoints with different heads only and are trained on the same data, leading to nearly identical behaviors that provide no corrective benefit. To tackle this critical challenge, we propose a simple yet effective non-overlapping data partitioning technique, in which each critic is trained via the subset formed by uniformly extracting responses from each prompt without overlap. This design encourages diversity among small critics and mitigates the risk of perception asynchrony among critics. Leveraging our ensemble-based value correction, small critics can provide reliable guidance to large policies despite their limited expressivity (Tint et al., 2024) (Figure 1(a)). Surprisingly, we find that double critic could be the sweet spot between correction capability and efficiency, it yields a qualitative leap in evaluation reliability while incurring the minimal redundancy needed for bias reduction. More critics increase the computation without proportional gains. Empirically, we demonstrate that two Qwen3-1.7b-Base critics robustly guide a larger policy, e.g., Qwen3-14b-Base, reducing critic over-parameterization while outperforming symmetric PPO under off-policy setting (Figure 1(b)). Notably, asymmetric architecture reduces peak memory by 20%, and accelerates training by around 20 seconds per step (see Figure 1(c)).
We further discovered that the agreement and divergence patterns in value estimates between our double critics, measured by their standard deviation, provide a useful signal for refining the policy loss objective. Value-estimation heterogeneity reflects both uncertainty and informativeness of the states. Leveraging this, we mask advantage values in states where critics strongly agree, reducing overfitting to low-quality samples and improving training stability. Conversely, we exploit divergence across critics by filtering out uncertain states from entropy regularization, since such states often correspond to low-probability continuations or spurious, reasoning-irrelevant patterns that inject noise into entropy measurements (Ahmed et al., 2019). Thus, restricting entropy regularization to high-confidence states promotes safer exploration and improves performance.
@@ 3.1 Towards Lightweight Value Estimation
Beyond explicitly increasing parameter differences through initialization, another promising approach is to provide differentiated optimization signals for each critic during training. Intuitively, training critics on non-overlapping subsets of data encourages them to learn from distinct trajectories and reward distributions, steering their updates in different directions and promoting functional diversity. However, in practice, randomly partitioning the training data can lead to asynchronous perception at the prompt level, where critics encounter inconsistent reasoning patterns from different questions. This imbalance increases the risk of overfitting to specific response types, resulting in unstable discrepancies in value estimates. In extreme cases, such divergence may cause policy collapse. To mitigate this, we uniformly divide the data into disjoint subsets at the prompt level, ensuring that each critic receives an equal share of responses within every prompt (or group). This design maintains perceptual synchrony across critics within each question while creating differentiated rewards and observations. Our ensemble critic system training process can then be formalized as:
@@ 3.2 Policy Loss Reconstruction
Beyond enabling robust value estimation, we conjecture that ensemble mini-critics can further enhance policy learning efficiency. Intuitively, the degree of agreement among critics’ value estimates for a given state can serve as a meaningful signal for policy optimization. This insight arises from our analysis of value fitting dynamics (Lee et al., 2021): when critics produce similar value estimates for a state $s_{i}$, it often indicates that $s_{i}$ is low-informative. Such states are frequently encountered across trajectories, and the rewards they yield exhibit low variance, causing critics to converge in their predictions, as visualized in Figure 4 (a). Analysis in Appendix D shows the positive correlation between value-std and the policy gradient, supporting the above speculation.
@@
Figure 5: Left: States with low value-std maintain low entropy (left box group), but states with low entropy may have a high value-std (right box group). Right: States with low entropy and states with low value-std show obvious differences. We sampled the $5\%$ lowest entropy states and the $5\%$ lowest value-std states at step 150.
Here, $\sigma_{t}=\text{std}\left(\{V(s_{t};\phi_{m})\}_{m=1}^{M}\right)$ denotes the agreement of value estimates for state $s_{t}$. Important sampling is defined as $\mathcal{IS}_{t}=\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{\text{old}}}(a_{t}|s_{t})}$. Figure 4(b) shows that, masking the advantages corresponding to the $20\%$ of states with high critic convergence, the policy exhibits stable learning dynamics even under high sample reuse (update-to-data ratio (UTD) = 4, i.e., each sample was used for training four times) and significantly improves sample efficiency. We further compared value-std (critic-side uncertainty) with entropy (policy-side uncertainty) (Wang et al., 2025a, Rahn et al., 2024, Cui et al., 2025b) by masking an equal fraction of states per step according to each metric. Figure 4(c) shows that value-std–based masking consistently delivers stronger learning benefits. This observation echoes classic RL findings (Osband et al., 2016), where ensemble-based value uncertainty acts as a proxy for learning dynamics. Figure 5 reveals that low value-std states consistently align with low entropy, suggesting that value-std is a precise uncertainty metric.
@@
When critics exhibit significant divergence in their evaluation of a state $s_{j}$, reflected in a high standard deviation, it may indicate that $s_{j}$ is reasoning-independent. For instance, different critics may encounter divergent reward distributions for trajectories passing through $s_{j}$, due to factors such as inference-irrelevant tokens or inherent semantic patterns in model generations. With a large $\lambda$, the dispersion in returns distribution propagates back to each state, amplifying disagreement among critics. In such cases, persistent exploration at $s_{j}$ is meaningless, as it does not correspond to an actionable decision state (Figure 6 (a)). To promote meaningful exploration while avoiding wasteful updates on noisy or non-decision states, we introduce a safe entropy regularization weighted by $\beta$. Specifically, we filter out states with high value standard deviation when computing entropy $\mathcal{H}$. Complete policy loss can be rewritten as:
@@
Figure 6 (b) shows that, unlike naive entropy loss, which can yield suboptimal learning, our entropy regularization mitigates entropy collapse and stabilizes policy learning, avoiding spurious exploration while guiding the policy toward better convergence with higher returns. We also compare filtering based on value-std versus entropy. As shown in Figure 6 (c), the overlap between the two sets is minimal. Even after filtering the top $40\%$ of hight value-std, policy entropy remains stable, while filtering the same fraction of high-entropy states causes entropy collapse. Statistical analysis of filtered tokens (Appendix E) further confirms that removed words are typically adverbs, interjections that are irrelevant to decision-making. Algorithm 1 summarizes the full execution process of AsyPPO.
@@ 4 Experiments
In §3, we described the architecture and training pipeline of AsyPPO and, through controlled ablation studies, demonstrated its efficacy on 4B and 8B LLMs (refer to Appendix B for detailed results). Building on that, this section examines AsyPPO more broadly through a suite of experiments. We organize the subsequent studies around three research questions: RQ1: Can AsyPPO and naive asymmetric PPO unlock general reasoning in larger LLM? RQ2: How sensitive is AsyPPO to the size and number of critics? RQ3: What setups are effective for advantage masking and entropy filtering?
@@ 4.1 Generalization to Large Models
For all algorithms, actors are initialized using Qwen3-14b-Base. Naive asymmetric PPO uses a single critic, i.e., Qwen3-1.7b-Base, Qwen3-4b-Base and Qwen3-8b-Base, and optimize with the vanilla PPO optimization objective. AsyPPO employs two mini-critics with advantage masking at $20\%$ and entropy filtering at $20\%$. We use the setting of GRPO recommended by Liu et al. (2025d). Full hyperparameter details are provided in Appendix C.2.
@@ 4.2 Ablation Study
The preceding results show that AsyPPO consistently enhances reasoning in base models across scales. We provide a module-wise analysis to characterize the algorithm from multiple perspectives.
@@
Figure 8 (a) shows a scaling-law–like trend: increasing critic size steadily raises the policy’s peak score. We recommend using the largest critic model that fits in GPU memory to maximize AsyPPO’s optimization capacity. However, we do not see comparable gains from increasing the number of critics: Figure 8 (b) shows that two mini-critics are sufficient for a clear step-change in performance. Varying the GRPO group size (trajectories per prompt) under our non-overlapping group setup while keeping other parameters at their defaults (Figure 8 (c)), and found 32 to be a robust setting. Comparing ensemble value aggregation (Figure 8 (d)), the mean of values outperforms min value, suggesting overestimation is not a dominant issue in RL4LLM.
@@ Acknowledgments
Appendix A Related work
Critic-based RL4LLM algorithms
On the other hand, value-based algorithm innovations are relatively few, Yuan et al. (2025b) argued that the decay factor is not well-suited for complex reasoning tasks that require long chains of thought (CoT). Yue et al. (2025), Zhu et al. (2025), Zhao et al. (2025) proposed novel mechanisms to enhance the robustness of the critic model when faced with noisy reward signals. Open-Reasoner-Zero (Hu et al., 2025) argues that, within this regime, vanilla PPO without KL regularization suffices to scale training stably. T-PPO (Fan et al., 2025) uses critic to enhance the stability of policy training in the long-tail asynchronous setting (Fu et al., 2025). Another similar research line to introduce critic-like models is done with the introduction of Implicit PRM (Yuan et al., 2025a). This approach is also able to provide token-level supervision for scalable RL training. PRIME (Cui et al., 2025a) adapted a specific reward model formulation to directly generate token-level rewards. However, current mainstream RL4LLM algorithms primarily emphasize critic-free optimization (Zhang et al., 2025). In this context, our research aim to underscore the importance of the critic in RL4LLM scenarios and try to address the deployment limitations associated with critics.
Asymmetric architecture.
In the realm of continuous deep RL, recent studies have investigated the potential of asymmetric network structures by reducing the capacity of the actor network. For example, Mastikhina et al. (2025), Mysore et al. (2021) suggest that the actor can function effectively with a significantly smaller capacity compared to the critic. Empirical evidence from Tan et al. (2022) supports this idea, demonstrating that sparsifying the policy network can enhance effective policy learning while significantly improving both inference and training speeds. Additionally, Liu et al. (2025a) found that pruning the actor network’s topology based on trial gradients can yield better performance. Similarly, Ma et al. (2025) revealed that even random pruning of the actor network can maintain performance within the SimBa network architecture (Lee et al., 2024). These contributions highlight the adaptability of RL in accommodating asymmetric designs, providing valuable insights for our research.
However, existing works primarily concentrate on reducing the actor’s size within simple network frameworks. In contrast, our paper pioneers the exploration of effectively guiding a small critic to inform a larger actor by optimizing the PPO algorithm within the RL4LLM scenario.
Appendix B The performance gain of AsyPPO on the small model strategy
Policy model Base model Symmetric PPO AsyPPO
Qwen3-4b-Base $30.5\%$ $47.3\%$ $53.1\%\;\scriptstyle{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+6.1\%}$
Qwen3-8b-Base $31.7\%$ $50.6\%$ $53.8\%\;\scriptstyle{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}+3.2\%}$
Table 1: Peak accuracy comparison of Symmetric PPO and AsyPPO under high data reuse setting (UTD=4) over six benchmarks. Score calculation same as Figure 1 (b). Purple score denotes the improvement compare to Symmetric PPO.
We set both the classic symmetrical PPO and our AsyPPO to the optimal Settings. AsyPPO uniformly initializes mini-critics using the Qwen3-1.7b-Base model. AsyPPO employs two mini-critics with advantage masking at 20%. And use the open source hard training dataset in (Liu et al., 2025d), which is selected from DeepMath-103k (He et al., 2025) with sampling probability proportional to each entry’s assigned difficulty level. We report the average@4 across six challenging benchmarks, i.e., MATH-500, OlympiadBench, MinervaMath, and AMC 2023, AIME 2025, AIME 2024.
Overall, Table 1 shows that AsyPPO effectively enhances the reasoning capabilities of two small models of different sizes, achieving respective improvements of 22.6% and 22.1% over their original performance. Compared to symmetric PPO, our algorithm delivers gains of 6.1% and 3.2%, while maintaining lightweight deployment. Upon analyzing specific benchmarks, our approach demonstrates notable advancements. For instance, on AIME 2025, we observed respective increases of approximately 4% (4B) and 6% (8B) compared to symmetric PPO. Similarly, on MATH-500, the improvements were around 3% (4B) and 2% (8B), and on MinervaMath, the gains were approximately 2% (4B) and 4% (8B). In the remaining three tasks, our method maintained performance levels comparable to those of symmetric PPO.
Appendix C Detailed Experimental Setup
C.1 Plot setup

### v2

@@ Abstract
Figure 1: (Left): Learnable critics naturally enhance policy stability through fine-grained value estimation and yield continuous gains as training progresses. Off-policy ratio=8, average@4 of 6 benchmarks, i.e., AIME 24, AIME 25, MATH-500, OlympiadBench, MinervaMath, and AMC 2023. (Right): AsyPPO restores the critic’s role in PPO while remaining lightweight and stable under LLM-scale training. The average clock time of training and the peak GPU memory usage of AsyPPO are significantly lower than those of the classic PPO, remain at the GRPO level.
@@ 1 Introduction
Faced with these challenges, the RL4LLM community has largely sidelined a key element of classical PPO – its critic. GRPO (He et al., 2025), and its variants, including GSPO (Zheng et al., 2025) in the Qwen series and DAPO (Yu et al., 2025), have achieved great success in replacing value functions with group sampling and average-advantage baselines for coarse-grained estimation of advantages. While effective, this paradigmatic shift abandons a key concept of RL: robust state value estimation can naturally mitigate training collapse caused by advantage bias (Wang et al., 2025b, Liu et al., 2024), especially under off-policy settings. This landscape motivates a fundamental reconsideration of architectural assumptions inherited from deep RL, prompting the following central question:
@@
Figure 2: Visual intuition behind AsyPPO. Small yet expressive critics can guide larger actors effectively when their representations are well-initialized. (a): A single critic struggles to align its value signal, leading to uncertain policy updates. Two small critics reach consensus (“This way”) and provide robust, low-variance guidance to the actor. (b): By leveraging representational priors and critic agreement, AsyPPO significantly outperforms classic PPO in terms of performance on various policy models, i.e., Qwem3-4b-Base and Qwem3-8b-Base. The Y-axis represents the improvement for the initial policy. The score calculation is the same as that in Figure 1.
To unlock the capabilities of the small critic, we consider critic ensembles to improve its value estimation and policy guidance. However, naive ensembles offer limited benefits for policy learning, as LLM critics start from identical pre-trained checkpoints with different heads only and are trained on the same data, leading to nearly identical behaviors that provide no corrective benefit. To tackle this critical challenge, we propose a simple yet effective non-overlapping data partitioning technique, in which each critic is trained via the subset formed by uniformly extracting responses from each prompt without overlap. This design encourages diversity among small critics and mitigates the risk of perception asynchrony among critics. Leveraging our ensemble-based value correction, small critics can provide reliable guidance to large policies despite their limited expressivity (Tint et al., 2024). Surprisingly, we find that double critic could be the sweet spot between correction capability and efficiency, it yields a qualitative leap in evaluation reliability while incurring the minimal redundancy needed for bias reduction. More critics increase the computation without proportional gains. Empirically, we demonstrate that two Qwen3-1.7b-Base critics robustly guide a larger policy, e.g., Qwen3-14b-Base, reducing critic over-parameterization while outperforming symmetric PPO under off-policy setting (Figure 1(a)). Notably, asymmetric architecture reduces peak memory by 20%, and accelerates training by around 20 seconds per step (see Figure 1(b)).
We further discovered that the agreement and divergence patterns in value estimates between our double critics, measured by their standard deviation, provide a useful signal for refining the policy loss objective. Value-estimation heterogeneity reflects both uncertainty and informativeness of the states. Leveraging this, we mask advantage values in states where critics strongly agree, reducing overfitting to low-quality samples and improving training stability. Conversely, we exploit divergence across critics by filtering out uncertain states from entropy regularization, since such states often correspond to low-probability continuations or spurious, reasoning-irrelevant patterns that inject noise into entropy measurements (Ahmed et al., 2019). Thus, restricting entropy regularization to high-confidence states promotes safer exploration and improves performance. With this refined loss objective, the policy further amplifies its learning efficiency ( Figure 2).
@@ 2 Preliminaries
3 Related work
Critic-based RL4LLM algorithms
On the other hand, value-based algorithm innovations are relatively few, Yuan et al. (2025b) argued that the decay factor is not well-suited for complex reasoning tasks that require long chains of thought (CoT). Yue et al. (2025), Zhu et al. (2025), Zhao et al. (2025) proposed novel mechanisms to enhance the robustness of the critic model when faced with noisy reward signals. Open-Reasoner-Zero (Hu et al., 2025) argues that vanilla PPO without KL regularization suffices to scale training stably. T-PPO (Fan et al., 2025) uses critic to enhance the stability of policy training in the long-tail asynchronous setting (Fu et al., 2025). Another similar research line to introduce critic-like models is done with the introduction of Implicit PRM (Yuan et al., 2025a). This approach is also able to provide token-level supervision for scalable RL training. PRIME (Cui et al., 2025a) adapted a specific reward model formulation to directly generate token-level rewards. However, current mainstream RL4LLM algorithms primarily emphasize critic-free optimization (Zhang et al., 2025). In this context, our research aim to underscore the importance of the critic in RL4LLM scenarios and try to address the deployment limitations associated with critics.
Asymmetric architecture.
In the realm of continuous deep RL, recent studies have investigated the potential of asymmetric network structures by reducing the capacity of the actor network. For example, Mastikhina et al. (2025), Mysore et al. (2021) suggest that the actor can function effectively with a significantly smaller capacity compared to the critic. Empirical evidence from Tan et al. (2022) supports this idea, demonstrating that sparsifying the policy network can enhance effective policy learning while significantly improving both inference and training speeds. Additionally, Liu et al. (2025a) found that pruning the actor network’s topology based on trial gradients can yield better performance. Similarly, Ma et al. (2025) revealed that even random pruning of the actor network can maintain performance within the SimBa network architecture (Lee et al., 2024). These contributions highlight the adaptability of RL in accommodating asymmetric designs, providing valuable insights for our research. However, existing works primarily concentrate on reducing the actor’s size within simple network frameworks. Our paper pioneers the exploration of effectively guiding small critics to inform larger actors by optimizing PPO within the RL4LLM scenario.
@@ 4.1 Towards Lightweight Value Estimation
Beyond explicitly increasing parameter differences through initialization, another promising approach is to provide differentiated optimization signals for each critic during training. Intuitively, training critics on non-overlapping subsets of data encourages them to learn from distinct trajectories and reward distributions, steering their updates in different directions and promoting functional diversity. However, in practice, randomly partitioning the training data can lead to asynchronous perception at the prompt level, where critics encounter inconsistent reasoning patterns from different questions. This imbalance increases the risk of overfitting to specific response types, resulting in unstable discrepancies in value estimates. In extreme cases, such divergence may cause policy collapse. To mitigate this, we uniformly divide the data into disjoint subsets at the prompt level, ensuring that each critic receives an equal share of responses within every prompt (or group). This design maintains perceptual synchrony across critics within each question while creating differentiated rewards and observations. Our ensemble critic training process can then be formalized as:
@@ 4.2 Policy Loss Reconstruction
Beyond enabling robust value estimation, we conjecture that ensemble mini-critics can further enhance policy learning efficiency. Intuitively, the degree of agreement among critics’ value estimates for a given state can serve as a meaningful signal for policy optimization. This insight arises from our analysis of value fitting dynamics (Lee et al., 2021): when critics produce similar value estimates for a state $s_{i}$, it often indicates that $s_{i}$ is low-informative. Such states are frequently encountered across trajectories, and the rewards they yield exhibit low variance, causing critics to converge in their predictions, as visualized in Figure 5 (a).
Figure 6: Left: At step 150, states with low value-std maintain low entropy (left box group), but states with low entropy may have a high value-std (right box group). Right: States with low entropy and states with low value-std show large differences.
Analysis in Appendix B shows the positive correlation between value-std and the policy gradient, supporting the above speculation.
@@
Here, $\sigma_{t}=\textit{std}\left(\{V(s_{t};\phi_{m})\}_{m=1}^{M}\right)$ denotes the agreement of value estimates for state $s_{t}$. Important sampling is defined as $\mathcal{IS}_{t}=\frac{\pi_{\theta}(a_{t}|s_{t})}{\pi_{\theta_{\text{old}}}(a_{t}|s_{t})}$. Figure 5(b) shows that, masking the advantages corresponding to the $20\%$ of states with high critic convergence, the policy exhibits stable learning dynamics even under high sample reuse (update-to-data ratio (UTD) = 4, i.e., each sample was used for training four times) and significantly improves sample efficiency. We further compared value-std (critic-side uncertainty) with entropy (policy-side uncertainty) (Wang et al., 2025a, Rahn et al., 2024, Cui et al., 2025b) by masking an equal fraction of states per step according to each metric. Figure 5(c) shows that value-std–based masking consistently delivers stronger learning benefits. This observation echoes classic RL findings (Osband et al., 2016), where ensemble-based value uncertainty acts as a proxy for learning dynamics. Figure 6 reveals that low value-std states consistently align with low entropy, suggesting that value-std is a precise uncertainty metric.
@@
When critics exhibit significant divergence in their evaluation of a state $s_{j}$, reflected in a high standard deviation, it may indicate that $s_{j}$ is reasoning-independent. For instance, different critics may encounter divergent reward distributions for trajectories passing through $s_{j}$, due to factors such as inference-irrelevant tokens or inherent semantic patterns in model generations. With a large $\lambda$, the dispersion in returns distribution propagates back to each state, amplifying disagreement among critics. In such cases, persistent exploration at $s_{j}$ is meaningless, as it does not correspond to an actionable decision state (Figure 7 (a)). To promote meaningful exploration while avoiding wasteful updates on noisy or non-decision states, we introduce a safe entropy regularization weighted by $\beta$. Specifically, we filter out states with high value std deviation when computing entropy $\mathcal{H}$. Complete policy loss is rewritten as:
@@
Figure 7 (b) shows that, unlike naive entropy loss, which can yield suboptimal learning, our entropy regularization mitigates entropy collapse and stabilizes policy learning, avoiding spurious exploration while guiding the policy toward better convergence with higher returns. We also compare filtering based on value-std versus entropy. As shown in Figure 7 (c), the overlap between the two sets is minimal. Even after filtering the top $40\%$ of hight value-std, policy entropy remains stable, while filtering the same fraction of high-entropy states causes entropy collapse. Statistical analysis of filtered tokens (Appendix C) confirms that removed words are typically adverbs, interjections that are irrelevant to decision-making. Algorithm 1 shows the pipeline of AsyPPO.
@@ 5 Experiments
In §4, we demonstrated the efficacy of AsyPPO on 4B and 8B LLMs (please refer to Figure 2(b) for results) through diversified experiments. This section examines AsyPPO more broadly through a suite of experiments. We organize the subsequent studies around three research questions: RQ1: Can AsyPPO and naive asymmetric PPO unlock general reasoning in larger LLM? RQ2: How sensitive is AsyPPO to the size and number of critics? RQ3: What setups are effective for advantage masking and entropy filtering?
@@ 5.1 Generalization to Large Models
For all algorithms, actors are initialized using Qwen3-14b-Base. Naive asymmetric PPO uses a single critic, i.e., Qwen3-1.7b-Base, Qwen3-4b-Base and Qwen3-8b-Base, and optimize with the vanilla PPO optimization objective. AsyPPO employs two mini-critics with advantage masking at $20\%$ and entropy filtering at $20\%$. We use the setting of GRPO recommended by Liu et al. (2025d). Full hyperparameter details are provided in Appendix A.3.
@@ 5.2 Ablation Study
The preceding results show that AsyPPO consistently enhances reasoning in LLMs across scales. We provide a module-wise analysis to characterize the algorithm from multiple perspectives.
@@
Figure 9 (a) shows a scaling-law–like trend: increasing critic size steadily raises the policy’s peak score. We recommend using the largest critic model that fits in GPU memory to maximize AsyPPO’s optimization capacity. However, we do not see comparable gains from increasing the number of critics: Figure 9 (b) shows that two mini-critics are sufficient for a clear step-change in performance. Varying the GRPO group size (trajectories per prompt) while keeping other parameters at their defaults (Figure 9 (c)), and found 32 to be a robust setting. Comparing ensemble value aggregation (Figure 9 (d)), the mean of values outperforms the min value, suggesting overestimation is not a dominant issue in RL4LLM.
@@ Acknowledgments
Appendix A Detailed Experimental Setup
A.1 Plot setup